# BEVTraj: Map-Free End-to-End Trajectory Prediction in Bird’s-Eye View with Deformable Attention and Sparse Goal Proposals

### v1

@@ III-B Iterative Deformable Decoder
The Iterative Deformable Decoder predicts and refines the multimodal future trajectory of the target agent based on $\mathbf{B}$ and $\mathbf{C}$, as illustrated in Fig. 4. It comprises three sub-modules: Sparse Goal Candidate Proposal, Initial Trajectory Prediction, and Iterative Trajectory Refinement. Each module employs deformable attention [32] to efficiently process the dense BEV representation. For simplicity, residual connections, normalization layers, feedforward networks, and some MLP components within the decoder are omitted in the following descriptions.
@@ III-B1 Sparse Goal Candidate Proposal
As illustrated in Fig.5, DenseTNT [20] produces dense goal candidates along lanes, while Motion TRansformer (MTR) [15] predefines intention points via k-means clustering on training data. We argue that both approaches require many goal candidates because they do not account for the dynamic state of the target agent or structural constraints when generating goals. To address this, we propose the Sparse Goal Candidate Proposal (SGCP) module, which predicts a sparse set of $\mathbf{K}$ goal candidates—significantly fewer than in prior methods—conditioned on the target agent’s dynamic state $\mathbf{s}\in\mathbb{R}^{t\times K_{attr}}$ and the BEV feature map $\mathbf{B}$.
[…]
Each block in SGCP follows this pipeline and can be stacked multiple times, progressively refining both $\mathbb{Q_{c}}$ and the predicted goal candidates.
Notably, when performing deformable cross-attention with $\mathbf{B}$, both the reference points and offsets are derived from the key—namely, the BEV feature map $\mathbf{B}$—rather than the query. Since $\mathbf{P}$ does not encode any road structure information, deriving sampling locations from $\mathbf{B}$ is more appropriate for identifying navigable regions.
Another key consideration in the design of SGCP is the coordinate system in which each operation is performed. Since $\mathbf{B}$ is defined in the ego-centric coordinate frame, deformable cross-attention must be performed in the ego space, requiring the dynamic state $\mathbf{s}$ to be represented in the same coordinate frame. In contrast, goal candidates and the corresponding initial trajectories of the target agent are predicted in the target-centric coordinate frame, meaning that mode self-attention must also operate in this space, with $\mathbf{s}$ accordingly transformed into the target agent’s coordinate frame. To ensure consistency in coordinate usage throughout the decoder, we apply deformable cross-attention prior to mode self-attention. More importantly, this ordering ensures that the $\mathbf{Q_{m}}$ contains rich contextual information, which is critical both for facilitating information exchange between $\mathbf{Q_{m}}$ during mode self-attention and for assigning distinct spatial roles to each mode query.
@@ III-B3 Iterative Trajectory Refinement
The ITR module operates as follows: it first applies temporal self-attention and mode self-attention to the decoder embedding output from ITP, denoted as $\in\mathbb{R}^{K\times T\times D}$. . It then performs cross-attention and deformable cross-attention with the scene context feature $\mathbf{C}$, where offsets are derived from the query embeddings, as in ITP. Finally, refinement offsets ($\Delta x,\Delta y$) are computed to update the trajectory predictions from the previous refinement block.
@@ III-C Losses
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{goal}}+\mathcal{L}_{\text{disp}}+\mathcal{L}_{\text{dense}}+\mathcal{L}_{\text{multi}}$ (1)
@@ III-C3 Dense Trajectory Loss ($\mathcal{L}_{\text{dense}}$)
$\mathcal{L}_{\text{dense}}=\lVert\hat{Y}-Y\rVert_{1}$
@@ III-C4 Multimodal Loss ($\mathcal{L}_{\text{multi}}$)
$\mathcal{L}_{\text{multi}}=\mathcal{L}_{\text{nll}}+\mathcal{L}_{\text{kl}}+\mathcal{L}_{\text{ent}}+\mathcal{L}_{\text{aux}}$ (4)
@@ IV Experiments
IV-A Datasets and Evaluation Metrics
IV-A1 Datasets
@@
Unlike nuScenes, which provides predefined target agents, Argoverse 2 Sensor does not specify prediction targets. Therefore, we extract 28k training samples and 6k validation samples by identifying suitable target agents based on object type, motion, trajectory validity, and proximity to the ego-vehicle. Each sample consists of 2 seconds of past and 6 seconds of future trajectory, sampled at 10 Hz. This setting is aligned with the nuScenes-based setup to ensure fair comparison.
IV-A2 Evaluation Metrics
To assess performance, The following evaluation metrics were adopted:
  • minADE (Minimum Average Displacement Error): The average Euclidean distance between the predicted trajectory and the ground-truth trajectory, computed over all future time steps. The minimum value among multiple predicted trajectories is used for evaluation, reflecting the best matching mode.
  • minFDE (Minimum Final Displacement Error): The Euclidean distance between the predicted final position and the ground-truth final position. Similar to minADE, the minimum distance among all predicted trajectories is taken.
  • Miss Rate: The proportion of predicted trajectories whose final displacement error exceeds a predefined threshold. Lower miss rates indicate more accurate and reliable predictions.
@@ IV-B1 Implementation Details
The overall BEVTraj architecture follows an encoder-decoder structure. Each sub-module in the Scene Context Encoder is designed as follows: the Pre-Encoder employs a temporal-social attention mechanism with two stacked layers, while the BDA module uses 256 BEV aggregate queries and applies deformable aggregation through three stacked layers. The local self-attention component consists of six layers in total.
The entire architecture uses a channel dimension of 256 for each token representation (e.g., content queries, positional queries) across all attention modules.
@@ IV-B2 Training Details
Prior to training for the trajectory prediction task, we pre-train the Sensor Encoder on a BEV map segmentation task to enhance its ability to capture static scene elements. The model is implemented in PyTorch and optimized using AdamW. The initial learning rate is set to $1\times 10^{-4}$ and decayed by a factor of 0.4 every 5 epochs, with a weight decay of 0.01. Dropout and layer normalization are applied to decrease overfitting. We train the model for 30 epochs with a batch size of 8 using four NVIDIA RTX 4090 GPUs. In order to jointly optimize different aspects of trajectory prediction,
our model is trained end-to-end using multiple loss functions shown in Eq.1
@@ IV-C Quantitative Results
To the best of our knowledge, this study is the first to attempt trajectory prediction using only raw sensor data without relying on pre-built HD maps. In this work, we aim to compare the performance of our proposed method with approaches that utilize HD maps. Unlike pre-built HD maps, sensor-based methods are inherently limited by the range of sensor data. For example, BEVFusion [6] utilizes a coverage of up to 50m, and many BEV models adopt the same data range [29]. As illustrated in Fig.6, while pre-built HD maps can provide access to static information regardless of the vehicle’s location—often using data within a ±100 m range—real-time HD maps are constrained by sensor range and are typically generated using information within ±50 m of the ego vehicle. Therefore, to ensure a fair comparison, we evaluate performance using HD map content limited to a ±50 m range around the ego vehicle.
@@
Tables I and II present a performance comparison between BEVTraj using 50m-range sensor data and prior approaches using 50 m-range of local HD maps. BEVTraj achieves strong results in minADE10, minFDE10, and Miss Rate (Top-10, 2m), while delivering comparable performance across other evaluation metrics. Notably, it excels in Miss Rate, which may be attributed to the presence of multiple samples where BEVTraj generates more plausible trajectories by leveraging information available exclusively from raw sensor data.
Unlike local HD map-based methods—which require expensive and labor-intensive construction—real-time encoding from raw sensor inputs introduces inherent challenges such as calibration errors, sensor noise, and information loss due to occlusion. Moreover, raw sensor data lacks explicit semantic and topological structure, including lane connectivity and road boundaries, which are readily accessible in pre-built HD maps. Nonetheless, despite these limitations, BEVTraj proves to be a competitive alternative under realistic conditions.
@@
Autobot 1.1417 0.6560 5.8954 1.3956 0.1762
MTR 0.7880 0.6799 4.6435 1.7082 0.2821
Wayformer 0.8191 0.5583 4.7608 1.3690 0.1837
DeMo 0.9000 0.6524 5.1404 1.3562 0.1776
BEVTraj (Ours) 0.9820 0.6249 5.2608 1.5832 0.1896
@@ IV-D Qualitative Results
To gain deeper insights into the prediction behavior of BEVTraj beyond quantitative metrics, we perform a qualitative study across diverse driving scenarios. This analysis aims to evaluate how effectively the model captures scene context and motion intent from raw sensor inputs.
As illustrated in Fig. 7, BEVTraj produces plausible future trajectories even under challenging conditions such as sharp curves and occluded intersections. While baseline models often generate trajectories that deviate into incorrect lanes, BEVTraj consistently preserves alignment with the correct driving lane. This suggests that raw sensor data may offer additional visual cues—such as road surface markings, physical barriers, and spatial continuity of free space—that are not explicitly represented in HD maps, which are typically constrained to predefined classes such as dividers and boundaries. We hypothesize that these fine-grained cues enable BEVTraj to more accurately infer drivable areas, thereby contributing to its superior performance in Miss Rate.
Although vectorized HD maps are compact and structurally efficient, they inherently abstract away the dense spatial details present in raw sensor data. In contrast, the BEV representation employed in BEVTraj preserves a spatial grid structure, which likely imposes stronger geometric constraints on predicted trajectories. This property appears particularly advantageous in scenarios involving sharp turns or lane merges, where precise alignment with road geometry is essential. In such cases, baseline models relying on vectorized maps often exhibit unstable lane-following behavior, whereas BEVTraj maintains more consistent and reliable performance.
This discrepancy may stem from architectural differences in the way each model encodes spatial information. For HD map-based models, the abstraction process may induce unstable activation distributions when encountering complex road geometries. In contrast, BEVTraj integrates explicit positional cues via deformable attention in a more granular and localized manner: spatial weights are not shared globally but are individually learned for each attention query.
In particular, the Iterative Deformable Decoder processes positional information independently for each predicted waypoint at every future timestamp. This design allows the model to adapt flexibly to local geometry, as each point along the predicted trajectory can selectively attend to spatial regions most relevant to its own position. Such per-timestep localization is especially advantageous in scenarios involving sharp curves or nonlinear motion, where uniform spatial treatment may degrade prediction accuracy.
Notably, BEVTraj succeeds in generating trajectories that align with implicit lane geometry and obstacle boundaries, even in the absence of explicit HD map information. These findings underscore the advantage of directly modeling dense features from raw sensor data, particularly in spatially complex environments where vectorized map representations may fall short.
@@
Figure 7: Qualitative comparison in diverse driving scenarios. BEVTraj shows consistently accurate, lane-aligned predictions under complex conditions such as sharp turns and occlusions.

### v2

@@ III-B Iterative Deformable Decoder
The Iterative Deformable Decoder predicts and refines the multimodal future trajectory of the target agent based on $\mathbf{B}$ and $\mathbf{C}$, as illustrated in Fig. 4. It comprises three sub-modules: Sparse Goal Candidate Proposal, Initial Trajectory Prediction, and Iterative Trajectory Refinement. Each module employs deformable attention [9] to selectively and adaptively aggregate agent-centric context from the dense BEV representation. For simplicity, residual connections, normalization layers, feedforward networks, and some MLP components within the decoder are omitted in the following descriptions.
@@ III-B1 Sparse Goal Candidate Proposal
As illustrated in Fig. 5, many existing trajectory prediction methods rely on generating a dense set of goal candidates, followed by trajectory exploration conditioned on these candidates. DenseTNT [53] samples goals densely along lanes, while Motion Transformer (MTR) [46] predefines intention points via k-means clustering on training data. Although such dense or anchor-based designs can alleviate mode collapse through brute-force diversity, they are inherently inflexible: most goal candidates are predefined and fail to adapt to scene-specific structures or the dynamic state of the target agent, and may even fall outside physically plausible or navigable regions. To overcome this limitation, we propose the Sparse Goal Candidate Proposal (SGCP) module, which directly predicts a small set of sample-adaptive, realistic, and reliable goal candidates. By explicitly conditioning goal generation on both the target agent’s dynamic state $\mathbf{s}$ and the BEV feature map $\mathbf{B}$, SGCP maintains strong prediction performance with only a limited number of sparse goals.
We next describe the architectural design and operational pipeline of SGCP in detail. SGCP adopts learnable seed parameters $\mathbf{P}\in\mathbb{R}^{K\times D}$ to represent $K$ latent behavior modes, where each seed functions as a mode query responsible for proposing one goal hypothesis. Unlike anchor-based approaches that rely on predefined spatial priors, these queries are learned end-to-end and must themselves maintain sufficient diversity. To incorporate the target agent’s motion history as an informative positional prior and to prevent different queries from collapsing into similar representations, the dynamic state $\mathbf{s}\in\mathbb{R}^{t\times K_{attr}}$ is first encoded by an MLP and injected into the mode queries through Feature-wise Linear Modulation (FiLM) [54]. Concretely, FiLM predicts query-specific scaling and shifting parameters conditioned on the history embedding and applies an affine transformation to each query independently. This lightweight conditioning modulates each query separately, enabling distinct, mode-dependent representations and explicitly mitigating mode collapse during early-stage goal generation.
To further incorporate scene context into the goal queries, we employ the BEV Deformable Aggregation (BDA) module, which is also used in the Scene Context Encoder, to selectively aggregate BEV features relevant to the target agent’s goal prediction. Specifically, decoder-side BDA aggregates features using target-aware reference points generated from the target agent’s motion history, collecting scene information from spatial regions most likely to influence the predicted goals. BDA also addresses a fundamental spatial misalignment issue: BEV features are defined in the ego-centric coordinate frame, whereas trajectories are predicted in the target-agent-centric frame. To resolve this discrepancy, the spatial coordinates of the aggregated BEV tokens are explicitly transformed into the target-agent-centric frame and used as positional keys for the subsequent cross-attention, producing mode-specific features $\mathbf{Q_{m}}\in\mathbb{R}^{K\times D}$.
The resulting mode features are further refined by a mode self-attention layer that models interactions and competition among the $K$ modes, yielding content queries $\mathbf{Q_{c}}\in\mathbb{R}^{K\times D}$. Placing this self-attention after the BEV cross-attention ensures that $\mathbf{Q_{m}}$ already encodes rich scene context, enabling meaningful information exchange across modes and promoting distinct spatial roles for each query. A lightweight MLP head then directly regresses $\mathbf{K}$ goal coordinates $\in\mathbb{R}^{K\times 2}$ in the target-centric space, generating a sparse set of realistic and behaviorally distinct goal candidates.
Beyond the above operational pipeline, we further highlight several architectural design choices that make SGCP efficient and robust, among which a key factor is the explicit decomposition of goal reasoning into two sequential stages. Decoder-side BDA first performs target-aware aggregation over the BEV feature map using a relatively large number of reference tokens (e.g., 256), forming a dense pool of goal-relevant scene features and providing broad spatial coverage (spatial exploration). A subsequent cross-attention then employs only a small set of mode queries to selectively attend to these aggregated tokens and refine them into sparse goal hypotheses (goal selection). This separation enables sufficient exploration while keeping the number of prediction modes compact. In contrast, methods such as MTR rely solely on mode queries for both exploration and selection, causing spatial coverage to scale with the number of modes and thus requiring many queries to capture diverse futures. By delegating dense exploration to BDA and reserving the queries for structured selection, SGCP achieves stable optimization and behaviorally diverse predictions with only a small set of realistic goal candidates.
In addition, the decoder-side BDA predicts reference points by encoding the target agent’s past trajectory under both target-centric and ego-centric coordinate systems to provide complementary motion cues for BEV aggregation. The target-centric representation offers normalized relative dynamics, while the ego-centric one maintains direct alignment with the BEV feature space. Combining both yields more reliable spatial anchoring and more effective feature aggregation.
@@ III-B3 Iterative Trajectory Refinement
The ITR module operates as follows: it first applies temporal self-attention and mode self-attention to the decoder embedding output from ITP, denoted as $\in\mathbb{R}^{K\times T\times D}$. It then performs cross-attention and deformable cross-attention with the scene context feature $\mathbf{C}$, where offsets are derived from the query embeddings, as in ITP. Finally, refinement offsets ($\Delta x,\Delta y$) are computed to update the trajectory predictions from the previous refinement block.
@@ III-C Losses
$\mathcal{L}=\lambda_{\text{goal}}\mathcal{L}_{\text{goal}}+\lambda_{\text{disp}}\mathcal{L}_{\text{disp}}+\lambda_{\text{dense}}\mathcal{L}_{\text{dense}}+\lambda_{\text{multi}}\mathcal{L}_{\text{multi}}$ (1)
@@ III-C3 Dense Trajectory Loss ($\mathcal{L}_{\text{dense}}$)
$\mathcal{L}_{\text{dense}}=\lVert\hat{Y}-Y\rVert_{1}$ (4)
@@ III-C4 Multimodal Loss ($\mathcal{L}_{\text{multi}}$)
$\mathcal{L}_{\text{multi}}=\lambda_{\text{nll}}\mathcal{L}_{\text{nll}}+\lambda_{\text{kl}}\mathcal{L}_{\text{kl}}+\lambda_{\text{ent}}\mathcal{L}_{\text{ent}}+\lambda_{\text{aux}}\mathcal{L}_{\text{aux}}$ (5)
@@
All loss weights are fixed across experiments. Specifically, we set $\lambda_{\text{goal}}=15$, $\lambda_{\text{disp}}=5$, $\lambda_{\text{dense}}=1$, $\lambda_{\text{multi}}=1$, $\lambda_{\text{nll}}=1$, $\lambda_{\text{ent}}=10$, $\lambda_{\text{KL}}=10$, and $\lambda_{\text{aux}}=100$.
@@ IV Experiments
IV-A Datasets
IV-A1 nuScenes
@@ IV-A2 Argoverse 2 Sensor dataset
Unlike nuScenes, which provides predefined target agents, Argoverse 2 Sensor does not specify prediction targets. Therefore, we extract 35k training samples and 7k validation samples by identifying suitable target agents based on object type, motion, trajectory validity, and proximity to the ego-vehicle. Each sample consists of 2 seconds of past and 6 seconds of future trajectory, sampled at 10 Hz. This setting is aligned with the nuScenes-based setup to ensure fair comparison.
@@ IV-B1 Implementation Details
The overall BEVTraj architecture follows an encoder-decoder structure. Each sub-module in the Scene Context Encoder is designed as follows: the Pre-Encoder employs a temporal-social attention mechanism with two stacked layers, while the BDA module uses 256 BEV aggregate queries and applies deformable aggregation through three stacked layers. The local self-attention component consists of six layers in total. The entire architecture uses a channel dimension of 256 for each token representation (e.g., content queries, positional queries) across all attention modules.
@@ IV-B2 Training Details
[…]
@@ IV-C Quantitative Results
IV-C1 Comparison with HD Map-Based Methods
@@
To the best of our knowledge, this study is the first to perform trajectory prediction using only raw sensor data without pre-built HD maps. Sensor-based perception is inherently limited to local, ego-centric observations, typically within 50 m in BEV-based pipelines such as BEVFusion [48] and BEVerse [61]. As illustrated in Fig. 6, real-time maps constructed from raw sensor data are naturally centered on the ego vehicle. Therefore, for a fair comparison, we restrict the HD map input of the map-based baseline methods to a ±50 m range centered at the ego vehicle when conducting our evaluation. Following the official nuScenes prediction benchmark, we report standard multimodal forecasting metrics including minADE, minFDE, and Miss Rate.
Tables I and II show that BEVTraj achieves competitive performance against HD map-based methods, with particularly strong results in Miss Rate. Despite lacking the explicit semantic and topological annotations of pre-built HD maps and being subject to sensing noise, BEVTraj remains competitive, demonstrating that reliable trajectory prediction is achievable without HD map dependency under realistic conditions.
@@
Autobot 1.3284 0.7430 6.9975 1.6861 0.2136
MTR 0.9243 0.7944 5.8264 1.9926 0.3361
Wayformer 1.0454 0.6721 6.2511 1.6775 0.2181
DeMo 0.8528 0.6426 5.2801 1.4964 0.1976
BEVTraj (Ours) 1.0412 0.6525 6.6665 1.6434 0.2144
TABLE III: The results of occupancy prediction evaluated over a $100\times 100$ m range.
Method IoU $\uparrow$
FIERY [62] 36.7
StretchBEV [63] 37.1
ST-P3 [64] 38.9
BEVerse [61] 40.9
PowerBEV [65] 39.3
UniAD [66] 40.2
FusionAD [67] 51.5
BEVTraj (GT-hist) 78.2
BEVTraj (E2E) 52.7
IV-C2 Scene-Level Evaluation via Occupancy Prediction
To complement per-agent trajectory metrics with a scene-level assessment, we additionally evaluate future occupancy prediction as a system-level proxy task. Unlike displacement-based errors that measure individual forecasts independently, occupancy forecasting directly reflects the spatial evolution of the entire scene and better captures transportation-system concerns such as space utilization, interaction feasibility, and potential conflict regions. For fair comparison with recent sensor-only end-to-end approaches that formulate motion prediction as BEV occupancy segmentation, we convert BEVTraj’s multi-agent trajectory outputs into occupancy grids through a deterministic post-processing step, without using any future ground-truth information. As summarized in Table III, BEVTraj achieves competitive performance under a realistic end-to-end setting where agent histories are inferred from detection and tracking results (E2E), and further demonstrates substantially stronger accuracy when ground-truth histories are provided (GT-hist). Despite not being explicitly trained with occupancy supervision, these results indicate that the proposed trajectory-centric modeling learns geometrically consistent and interaction-aware scene dynamics. Overall, the occupancy evaluation verifies that BEVTraj generalizes beyond agent-level forecasting and maintains coherent system-level scene prediction.
IV-C3 Robustness under Challenging Driving Conditions
To assess robustness under adverse driving conditions, we conduct a scenario-wise evaluation on the nuScenes validation set, including rain, night, construction, and heavy_traffic, where all subsets are consistently filtered using scenario metadata and evaluated with the same protocol and metrics as the full set (Table IV). We focus on the failure-oriented metric Miss Rate10, which more directly reflects collapse-like prediction behavior. BEVTraj maintains consistently low failure rates across all subsets (0.22–0.30), achieving particularly strong robustness at night and remaining stable in complex scenes such as construction and heavy_traffic, where occlusions and dense interactions are prevalent. By contrast, map-based approaches rely on pre-built HD maps and are therefore largely unaffected by such adverse perception conditions. Overall, these results demonstrate that the proposed map-free design provides reliable and stable trajectory predictions across diverse and challenging environments.
TABLE IV: Scenario-wise performance on the nuScenes validation set. We report results on four challenging conditions: rain, night, construction, and heavy_traffic.
Condition Method minADE5 minADE10 minFDE1 minFDE10 Miss Rate10
rain Autobot 1.9784 1.1499 8.7597 2.2758 0.3133
MTR 1.2967 1.0627 7.2870 2.3351 0.4357
Wayformer 1.3827 0.9661 8.0107 2.1527 0.3793
DeMo 1.2927 0.8345 6.4617 1.6804 0.2256
BEVTraj (Ours) 1.4626 0.9464 8.3420 2.0262 0.2995
night Autobot 1.9710 1.1554 8.9577 2.3311 0.3170
MTR 1.4276 1.1298 7.7650 2.4510 0.4313
Wayformer 1.3492 0.9828 7.7897 2.2732 0.3825
DeMo 1.2812 0.8654 6.2776 1.7833 0.2202
BEVTraj (Ours) 1.2870 0.9034 7.0542 1.9268 0.2201
construction Autobot 1.9572 1.1795 8.7640 2.4068 0.3431
MTR 1.2964 1.0554 7.1921 2.3023 0.4232
Wayformer 1.3059 0.9564 7.7272 2.1423 0.3540
DeMo 1.2952 0.8551 6.3589 1.7188 0.2371
BEVTraj (Ours) 1.4480 0.8965 8.4993 1.8684 0.2750
heavy_traffic Autobot 0.8487 0.7962 5.7484 2.0362 0.3208
MTR 0.9042 0.8449 3.6851 2.0185 0.5094
Wayformer 0.7669 0.6949 6.2221 2.0471 0.3774
DeMo 0.6723 0.5962 2.8915 1.4073 0.1887
BEVTraj (Ours) 0.8746 0.6621 5.6646 1.6686 0.3019
@@ IV-D Qualitative Results
To gain deeper insights into the prediction behavior of BEVTraj beyond quantitative metrics, we perform a qualitative study across diverse driving scenarios. This analysis aims to examine how effectively the model extracts motion-relevant scene context and surrounding agent cues from raw sensor inputs, and how these representations translate into consistent and plausible trajectory predictions.
IV-D1 Qualitative Comparison of Trajectory Predictions
As illustrated in Fig. 7, BEVTraj produces stable and lane-aligned future trajectories even under challenging conditions such as sharp curves and cluttered intersections. In contrast, baseline models frequently generate trajectories that drift into incorrect lanes or violate road boundaries, particularly in geometrically complex regions.
@@ IV-D1 Qualitative Comparison of Trajectory Predictions
Figure 7: Qualitative comparison in diverse driving scenarios. BEVTraj shows consistently accurate, lane-aligned predictions under complex conditions such as sharp turns and cluttered intersections, benefiting from directly leveraging raw sensor cues that are not available in map-based representations.
We attribute this difference primarily to the nature of the input representation. Vectorized HD maps provide compact and structured abstractions, but their coordinate-based representations are typically encoded directly by MLP-based networks, which may tend to weaken local spatial correlations and inevitably discard fine-grained spatial details present in raw sensor data. As a result, fine-grained cues such as road surface markings, free-space continuity, and subtle geometric boundaries are not explicitly preserved. In contrast, BEVTraj operates on a rasterized BEV representation constructed directly from sensors, which maintains continuous spatial structure and richer local geometry, allowing the model to better infer drivable areas and respect lane constraints.
These observations suggest that directly modeling dense sensor-derived features provides stronger geometric guidance for trajectory prediction, contributing to BEVTraj’s improved robustness in complex driving scenarios.
IV-D2 Visualization of BEV Feature Representations
To better understand why the BEV formulation provides stronger geometric guidance, we directly visualize the intermediate BEV features constructed from camera, LiDAR, and their fusion (Fig. 8). These visualizations reveal clear modality-specific characteristics. Camera-derived BEV features tend to emphasize semantic and contextual cues, while LiDAR-derived features exhibit sharper geometric structures corresponding to road boundaries, obstacles, and free-space layouts.
Figure 8: Qualitative visualization of the BEV representation prior to trajectory reasoning. (a) Fused BEV feature representation constructed from multi-view cameras and LiDAR, where the right side corresponds to the ego vehicle’s forward direction. (b) Corresponding multi-view camera images used jointly with LiDAR to build the fused BEV representation.
Figure 9: Per-channel visualization of the fused BEV feature. Different channels exhibit distinct activation patterns, reflecting heterogeneous semantic and geometric information preserved in the fused BEV representation.
The fused BEV feature is visualized via channel-wise max pooling, which aggregates heterogeneous activations into a single dense map and makes it difficult to interpret directly. A clearer picture emerges when we inspect the features on a per-channel basis (Fig. 9). Different channels exhibit markedly different activation patterns: some resemble geometry-dominant structures similar to LiDAR features, while others reflect semantic activations akin to camera-based features. This heterogeneous structure suggests that the fused BEV representation implicitly preserves complementary semantic and geometric information across channels, providing rich scene cues that can be selectively exploited for downstream trajectory reasoning.
IV-D3 Dense Scene-Level Motion Modeling
Beyond predicting the designated target agent, BEVTraj explicitly accounts for the future motions of surrounding agents through dense future prediction [46]. During scene context encoding, the model predicts dense multi-agent futures over the BEV grid, enabling scene-level motion understanding that captures how nearby agents are expected to evolve over time.
As shown in Fig. 10, the resulting predictions exhibit coherent and consistent multi-agent motion patterns across the scene. This dense consideration of surrounding traffic provides informative interaction cues and contributes to more stable and interaction-aware target-agent predictions.
Figure 10: Visualization of dense future prediction for surrounding agents.
Figure 11: Visualization of SGCP-proposed goals and their corresponding final trajectories. SGCP can generate a compact, spatially diversified set of goal candidates.
IV-D4 Sparse Goal Candidates and Decoded Trajectories
To illustrate how BEVTraj forms structured motion hypotheses prior to trajectory decoding, we visualize the outputs of the Sparse Goal Candidate Proposal (SGCP) module in Fig. 11. SGCP generates a compact and spatially diversified set of goal candidates that represent distinct high-level motion intents, enabling efficient hypothesis generation directly from BEV features without relying on predefined structures or hand-crafted sampling.
Each goal candidate is overlaid with its corresponding decoded trajectory, explicitly showing how sparse goals serve as anchors for multi-modal prediction. The decoder subsequently refines these anchors into temporally consistent motion forecasts, linking high-level goal reasoning with detailed trajectory generation.
IV-D5 Layer-wise Evolution of Deformable Attention Geometry
To illustrate how BEVTraj progressively refines future trajectories, we visualize the layer-wise evolution of reference points and sampling locations in the Iterative Trajectory Refinement (ITR) module in Fig. 12. Each layer updates future waypoint references, which subsequently guide deformable attention to aggregate scene features around the current motion hypotheses.
Figure 12: Visualization of iterative reference point refinement in the Iterative Trajectory Refinement (ITR) module. The top panel shows the six camera images. The subsequent panel visualizes the predicted future trajectory as deformable-attention reference points, together with the corresponding sampling locations, overlaid with the ground-truth future trajectory in the ego-vehicle–centered coordinate frame.
Across iterations, the sampling locations adaptively concentrate on regions that are most relevant to the predicted path, yielding a coarse-to-fine refinement process. This dynamic receptive field enables agent-centric and motion-aware feature aggregation, allowing the model to progressively improve trajectory accuracy without relying on explicit HD-map priors.
IV-D6 Robustness under Adverse Visibility Conditions
Finally, to complement the representation and mechanism-oriented analyses above, we evaluate the robustness of BEVTraj under adverse visibility conditions such as rain and night, where visibility is degraded by raindrops, reflections, and low illumination. Fig. 13 provides qualitative comparisons with multiple representative methods, offering a direct visual reference of trajectory forecasts under these challenging scenarios, particularly during geometry-sensitive maneuvers such as turning and lane-following. In the shown examples, BEVTraj maintains trajectories that remain well aligned with the local road geometry and closely follow the ground-truth motion, demonstrating stable lane-consistent behavior despite weakened visual cues. These results further confirm that BEVTraj produces reliable and usable predictions even under degraded-visibility conditions.
Figure 13: Qualitative comparison of trajectory predictions under adverse visibility in nuScenes. Each column presents one representative case (rain/night). The top row shows the front-view camera image, and the bottom row overlays multi-modal trajectory predictions from different methods in the BEV space together with the ground truth. For each method, we visualize the highest-score trajectory among its multiple candidates.